OpenStudy (anonymous):

A sphere is inscribed in a cylinder. The surface area of the sphere and the lateral area of the cylinder are equal. true or false

OpenStudy (valpey):

True. The surface area of the sphere is: \[4\pi r^2\] The lateral surface area of the cylinder is Circumference x 2 x radius: \[2\pi r * 2r = 4\pi r^2\]